Your are going to solve a 3-by-3 system of equations. You will be guided through it step-by- step, as this is a bit of an advanced problem.
Equation 1: x + y + z = 2
Equation 2: x + y - z = -2
Equation 3: 2x - y + z = 5
Step 1: Combine equation 1 and 2 to get an equation only in x and y. Call this equation 4.
Step 2: Combine equation 1 and 3 to get another equation only in x and y. Call this equation 5. You will need to slightly modify one of the equations to eliminate z.
Step 3: Now that you have two equations in terms of only x and y ( Equations 4 and 5), use any method of your choice to solve for x and y.
Step 4: Now that you have x and y, solve for z using any of the first three equations.

Found 2 solutions by josgarithmetic, MathTherapy:
Answer by josgarithmetic(39617)

The problems are not identical but the method could or should be similar - Row Reduction, matrices.

https://www.algebra.com/algebra/homework/Systems-of-equations/Systems-of-equations.faq.question.1042299.html

Answer by MathTherapy(10552)


Your are going to solve a 3-by-3 system of equations. You will be guided through it step-by- step, as this is a bit of an advanced problem.
Equation 1: x + y + z = 2
Equation 2: x + y - z = -2
Equation 3: 2x - y + z = 5
Step 1: Combine equation 1 and 2 to get an equation only in x and y. Call this equation 4.
Step 2: Combine equation 1 and 3 to get another equation only in x and y. Call this equation 5. You will need to slightly modify one of the equations to eliminate z.
Step 3: Now that you have two equations in terms of only x and y ( Equations 4 and 5), use any method of your choice to solve for x and y.
Step 4: Now that you have x and y, solve for z using any of the first three equations.
The steps are all there for you to solve the system. What's holding you back?